# On the Learnability of Watermarks for Language Models

### v2

@@ 5.3 Robustness to changes in decoding algorithm
KGW $\delta=2$ 6.9e-09 2.9e-09 1.2e-09 4.8e-10 1.9e-10 2.8e-11 3.7e-12 1.3e-12
Aar $k=2$ 2.2e-12 4.6e-13 8.5e-14 3.9e-14 5.2e-15 9.9e-17 9.9e-18 5.3e-18
KTH $s=1$ 1.0e-04 1.0e-04 1.0e-04 1.0e-04 1.0e-04 1.0e-04 1.0e-04 1.0e-04
Table 3: Watermark detection p-values of generations from logits-based watermark distilled Llama 2 7B models under various decoding algorithms and parameters. We observe small p-values in all settings, showing that weights-based watermarking is robust to changes in the decoding algorithm.
Whereas decoding-based watermarking relies on special decoding algorithms, weights-based watermarking generates watermarked text naturally, i.e., under standard decoding algorithms. In Table 3, we compute watermark detection p-values from weights-based watermarking under a variety of decoding algorithms and parameters: nucleus sampling (Holtzman et al., 2020) with different $p$, temperature-based sampling with different temperatures $t$, and greedy decoding ($t=0$).
We find that weights-based watermarking is robust to changes in the decoding algorithm, achieving consistently high detectability. We also observe that p-values decrease as $t$ or $p$ decreases, with greedy decoding obtaining the smallest p-values.
@@ 5.4 Robustness to text edits
We test the robustness of weights-based watermarking to edits by randomly corrupting generated text from logits-based watermark distilled Llama 2 7B models with varying proportions of tokens randomly edited. See Appendix G for full experimental details. As shown in Figure 4, all three watermarks are robust to moderate proportions of tokens edited, up to around 0.2-0.3. The watermark remains detectable with small p-values. At higher proportions of random edits, up to around 0.6, KTH is significantly more robust to edits than KGW and Aar, consistent with the findings of Kuditipudi et al. (2023).
Figure 3: Watermark detection p-values on generations from weights-based watermarking, corrupted with varying proportions of tokens randomly edited. The watermarks are robust to moderate amounts of corruption.
Figure 3: Watermark detection p-values on generations from weights-based watermarking, corrupted with varying proportions of tokens randomly edited. The watermarks are robust to moderate amounts of corruption.
Figure 4: Watermark detection p-values of generations from logits-based watermark distilled Llama 2 7B models after further fine-tuning on OpenWebText. The models’ weights-based watermarking is removed by fine-tuning.
6 Applications
6.1 Watermarking for open models
In §5.3, we showed that weights-based watermarking is robust to changes in the decoding algorithm. This is a necessary first step towards watermarking for open models, where users can run inference themselves. They may change the decoding algorithm, and the inference library they use may not enable decoding-based watermarking by default or implement it at all.
Robust watermarking for open models should also ideally be robust to fine-tuning, as users have the ability and desire to fine-tune open models for their own use cases. However, Figure 4 shows that weights-based watermarking obtained from watermark distillation is not robust to further fine-tuning on normal, non-watermarked text (see Appendix H for experimental details). We leave addressing this challenge and learning weights-based watermarking that is robust to fine-tuning to future work.
@@ 6.1 Watermarking for open models
6.2 Spoofing attacks
Abdelnabi & Fritz (2021) and Kuditipudi et al. (2023) consider the possibility of using watermark detection to attribute the provenance of generated text to a specific model, which could help policy enforcement and auditing of model providers. However, using watermarking for provenance attribution brings the risk of spoofing attacks: an adversary generates damaging text containing the watermark of a victim model, making it appear as if the victim model generated it, thus hurting the reputation of the victim model (Sadasivan et al., 2023). Sampling-based watermark distillation is applicable to the spoofing attack setting, as only black-box samples from the victim/teacher model are needed for distillation.
In this proof-of-concept experiment, we simulate a spoofing attack using a victim model of Llama 2-Chat 7B with the KGW decoding-based watermarking strategy ($\delta=2$). Llama 2-Chat 7B is trained for safety and tends to refuse harmful requests (Touvron et al., 2023). The goal of the spoofing attack is generate watermarked responses to harmful requests, damaging the victim model’s reputation for safety. We obtain an adversary model by performing sampling-based watermark distillation with Alpaca-7B (Taori et al., 2023) as the student and the victim model Llama 2-Chat 7B as the teacher. We query the victim model for watermarked samples, filter out refusals, then fine-tune the adversary model on those samples. See Appendix I.1 for full training details.
@@ 7 Related work
Text watermarking. Older works on textual watermarking edit pre-existing text to inject signals that can be statistically detected (Rizzo et al., 2019; Abdelnabi & Fritz, 2021; Yang et al., 2022), see Kamaruddin et al. (2018) for a survey. More specifically, we study decoding-based watermarking which modify decoding procedures to generate new watermarked text (Venugopal et al., 2011; Kirchenbauer et al., 2023a; Aaronson, 2023; Kuditipudi et al., 2023; Zhao et al., 2023a; Christ et al., 2023; Hu et al., 2023; Wu et al., 2023). Recent works have proposed various watermarking algorithms, including semantic watermarks (Fu et al., 2023; Hou et al., 2023; Liu et al., 2023b; Ren et al., 2023), multi-bit watermarking (Yoo et al., 2023; Wang et al., 2023), and public/private key watermarking (Liu et al., 2023a; Fairoze et al., 2023), but note that these are also all decoding-based. Decoding-based watermarking is not robust to changes in decoding algorithm, whereas we find that weights-based watermarking is.
Watermark attacks. Prior works have investigated attacks to remove the watermark from watermarked text, e.g., by paraphrasing, swapping tokens, etc. (Kirchenbauer et al., 2023b; Krishna et al., 2023; Sadasivan et al., 2023; Zhang et al., 2023). In addition, watermark spoofing attacks are where an adversary generates watermarked text that falsely appears as if it came from a victim model. Sadasivan et al. (2023) investigates some spoofing attacks, but they require knowledge of how the watermark signal works. Our spoofing attack using sampling-based watermark distillation does not requre such knowledge.
Watermarking for protection against model extraction. A number of prior works have studied watermarking for protection against model extraction/stealing attacks, where an adversary “steals” a model via API access by distilling on its outputs (He et al., 2022a; Zhao et al., 2022; He et al., 2022b; Zhao et al., 2023b). In this setting, the suspect model is queried multiple times on specially chosen inputs to detect whether the model was stolen. The watermark is specifically designed to be learnable by the adversary model via distillation. On the other hand, text watermarking enables detecting whether any text was model-generated from any input prompt, so the object of detection is different (model versus text). Existing decoding-based text watermarking strategies are designed to be subtle and imperceptible in order to minimally impact text quality. They are not specifically designed to be learnable, so their learnability must be determined via empirical investigation.
@@ 8 Conclusion
In this paper, we investigate the learnability of watermarks for language models. We find that logits-based and sampling-based watermark distillation are able to successfully learn weights-based watermarking that is robust to changes in decoding algorithm and text edits, although lower-distortion watermarks are harder and less sample efficient to learn. Our findings address a key technical challenge towards developing watermarking for open models and raise the possibility of watermark spoofing attacks against victim models.
Future work may explore improving robustness of weights-based watermarking to further fine-tuning, which would address another important challenge towards robust watermarking for open models. Future work may also more comprehensively study and evaluate spoofing attacks and potential defenses against spoofing attacks, which would have implications for whether watermarks should be used to assign provenance and blame to specific models.
@@ Acknowledgments
Appendix A Watermark distillation training details
A.1 Logits-based watermark distillation training details
@@
We train using a subset of OpenWebText for 5,000 steps with a batch size of 64, sequence length of 512 tokens, maximal learning rate of 1e-5, and cosine learning rate decay with a linear warmup for the first 500 steps, and the AdamW optimizer (Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Each training run took approximately 6 hours on 4 NVIDIA A100 80GB GPUs.
A.2 Sampling-based watermark distillation training details
@@
We fine-tune Pythia 1.4B on the watermarked samples for 1 epoch, roughly 8,000 steps, with a batch size of 64, sequence length of 256 tokens, maximal learning rate of 1e-5, and cosine learning rate decay with a linear warmup for the first 500 steps, and the AdamW optimizer (Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Each training run took approximately 3 hours on 1 NVIDIA A100 80GB GPU.
Appendix B Full sampling-based watermark distillation results table
Table 4 contains the full version of Table 2, including values for the seq-rep-3 metric, which was deferred here due to space constraints. We observe the same trends for seq-rep-3 as in logits-based watermark distillation (Table 1), where Aar increases repetition in all methods, especially for lower $k$, whereas KGW and KTH do not affect repetition.

### v3

@@ 5.1 Robustness to changes in decoding parameters
Logit-based watermark distillation
KGW $k=1,\delta=2$ 7e-09 3e-09 1e-09 7e-10 2e-10 1e-11 4e-12 1e-12
Aar $k=2$ 2e-12 3e-13 6e-14 3e-14 3e-15 6e-17 6e-18 5e-18
KTH $s=1$ 1e-04 1e-04 1e-04 1e-04 1e-04 1e-04 1e-04 1e-04
Sampling-based watermark distillation
KGW $k=1,\delta=2$ 1e-06 4e-07 2e-07 2e-07 4e-08 3e-09 3e-09 5e-09
Aar $k=2$ 1e-15 7e-16 1e-16 1e-17 1e-18 3e-21 1e-22 2e-22
KTH $s=1$ 1e-04 1e-04 1e-04 1e-04 1e-04 1e-04 1e-04 1e-04
Table 2: Watermark detection p-values of generations from logit-based and sampling-based watermark distilled Llama 2 7B models under various decoding parameters. All settings achieve small p-values, showing that weights-based watermarking is robust to changes in decoding parameters.
@@ 5.2 Robustness to text edits
We test the robustness of weights-based watermarking to edits by randomly corrupting generated text from the logit-based and sampling-based watermark distilled Llama 2 7B models with varying proportions of tokens randomly edited. See Appendix G for full experimental details. As shown in Figure 4, the detection p-values of all three watermark types are robust to moderate edit proportions, up to around 20–30%. At higher edit proportions, up to around 60–70%, KTH is significantly more robust to edits than KGW and Aar, consistent with the findings of Kuditipudi et al. (2023).
Figure 3: Watermark detection p-values of generations from weights-based watermarking, corrupted with varying proportions of tokens randomly edited. The watermarks are robust to moderate amounts of corruption.
Figure 4: Watermark detection p-values of generations from logit-based watermark distilled Llama 2 7B models after further fine-tuning on OpenWebText. The models’ weights-based watermarking is removed by fine-tuning.
6 Watermarking for open models
In §5, we showed that weights-based watermarking works under standard decoding algorithms and is robust to changes in decoding parameters. This is a necessary first step towards watermarking for open models, where users can run inference themselves. They may change the decoding algorithm, and the inference library they use may not enable decoding-based watermarking by default or implement it at all.
Robust watermarking for open models should also ideally be robust to fine-tuning, as users have the ability and desire to fine-tune open models. Ideally, this fine-tuning should not remove watermarking capabilities, either intentionally or unintentionally. However, Figure 4 shows that weights-based watermarking obtained from watermark distillation is not robust to further fine-tuning on normal, non-watermarked text (see Appendix H for experimental details). We leave addressing this challenge and learning weights-based watermarking that is robust to fine-tuning to future work.
@@ 6 Watermarking for open models
7 Spoofing attacks
One proposed use case of watermark detection is to attribute the provenance of generated text to a specific model, which could help policy enforcement and auditing of model providers (Abdelnabi & Fritz, 2021; Kuditipudi et al., 2023). However, using watermarking for provenance attribution brings the risk of spoofing attacks: an adversary generates damaging text containing the watermark of a victim model, making it appear as if the victim model generated it, thus hurting the reputation of the victim model (Sadasivan et al., 2023). Sampling-based watermark distillation is applicable to the spoofing setting, as it only requires generated samples from the victim/teacher model.
In this proof-of-concept experiment, we simulate a spoofing attack using a victim model of Llama 2-Chat 7B with KGW decoding-based watermarking ($k=1,\gamma=0.25,\delta=2$). Llama 2-Chat 7B is trained for safety and tends to refuse harmful requests (Touvron et al., 2023). The goal of the spoofing attack is to generate watermarked responses to harmful requests, damaging the victim model’s reputation for safety. We obtain an adversary model by performing sampling-based watermark distillation with Alpaca-7B (Taori et al., 2023) as the student and the Llama 2-Chat 7B victim model as the teacher. We query the victim model for watermarked samples, filter out refusals, then fine-tune the adversary model on those samples. See Appendix I.1 for full experimental details.
@@ 8 Related work
Text watermarking. Older works on text watermarking edit pre-existing text to inject signals that can be statistically detected (Rizzo et al., 2019; Abdelnabi & Fritz, 2021; Yang et al., 2022), see Kamaruddin et al. (2018) for a survey. Recently, many works have studied decoding-based watermarking, which modifies decoding procedures to generate new watermarked text (Venugopal et al., 2011; Kirchenbauer et al., 2023a; Aaronson, 2023; Kuditipudi et al., 2023; Zhao et al., 2023a; Christ et al., 2023; Hu et al., 2023; Wu et al., 2023; Huang et al., 2023; Zhao et al., 2024). Various classes of decoding-based watermarking methods have been proposed, e.g., semantic watermarks (Fu et al., 2023; Hou et al., 2023; Liu et al., 2023b; Ren et al., 2023), multi-bit watermarking (Yoo et al., 2023; Wang et al., 2023; Qu et al., 2024; Boroujeny et al., 2024), and public/private key watermarking (Liu et al., 2023a; Fairoze et al., 2023). See (Liu et al., 2023c) for a survey of text watermarking. Sander et al. (2024) find that it is possible to detect if a model’s training data contained watermarked text.
Watermark attacks. Recent works have investigated attacks to remove the watermark from watermarked text, using methods such as paraphrasing, swapping tokens, etc. (Kirchenbauer et al., 2023b; Krishna et al., 2023; Sadasivan et al., 2023; Zhang et al., 2023; Pang et al., 2024; Jovanović et al., 2024). In addition, watermark spoofing attacks are where an adversary produces text that is falsely detected as watermarked and generated by a victim model. Sadasivan et al. (2023) and Jovanović et al. (2024) spoof the KGW watermark by exploiting its green list bias, and Pang et al. (2024) demonstrate spoofing attacks by exploiting watermark robustness and public detection APIs. In our work, we show that sampling-based watermark distillation can enable spoofing attacks.
API watermarking for protection against model extraction. Prior works have studied API watermarking for protection against model extraction attacks, where an adversary imitates or reconstructs a victim model by distilling on its API outputs (He et al., 2022a; Zhao et al., 2022; He et al., 2022b; Zhao et al., 2023b). In API watermarking, a watermark signal is injected into the victim’s API outputs, making it possible to detect if a suspect model was distilled from the victim API. In contrast, text watermarking enables detecting whether a given text was model-generated.
@@ 9 Conclusion
In this paper, we investigate the learnability of watermarks for language models. Using logit-based and sampling-based watermark distillation, we find that models can learn to naturally generate watermarked text using standard decoding algorithms, although lower-distortion watermarks are harder and less sample efficient to learn. Our findings address a key technical challenge towards developing watermarking for open models and raise the possibility of watermark spoofing attacks.
Future work may explore improving the robustness of weights-based watermarking to further fine-tuning, which would address another important challenge towards robust watermarking for open models. Future work may also more comprehensively study and evaluate spoofing attacks and potential defenses against spoofing attacks, which would have implications for whether watermarks should be used to assign provenance and blame to a specific model.
Ethics statement
In this paper, we find that sampling-based watermark distillation can potentially be used to carry out harmful watermark spoofing attacks. This may appear to be a potentially harmful insight that weakens watermarking by undermining its ability to identify the provenance of text. However, we believe that public knowledge of spoofing attacks and the limitations of watermarking is important. This way, the public knows not to trust watermarking for reliably attributing provenance or blame to a specific model. Then, if watermark detection is not used to prove that a text was generated by a specific model, spoofing attacks will cause significantly less harm, if any at all. Watermarking can still be used to statistically detect LM-generated text, which can be used for tasks such as finding infractions of policies on language model usage.
Reproducibility statement
For the main results, we describe our experimental setup in §4, including training details, datasets used, and evaluation procedure. For all other experiments and results, we describe full experimental details in the appendix. The exact sections in the appendix are mentioned in the main paper where relevant. In addition, we release code and scripts to reproduce experiments at https://github.com/chenchenygu/watermark-learnability, along with trained model weights.
@@ Acknowledgments
Appendix A Additional details on watermarking strategies
In this section, we provide additional details and formal definitions for the KGW, Aar, and KTH watermarking strategies, using the definitions of decoding-based watermarking strategies (equation 1) and watermark detection (equation 2).
A.1 KGW
Formally, the KGW (Kirchenbauer et al., 2023a) decoding-based watermarking strategy can be defined as
$f_{\text{w}}^{\text{KGW}}\left({\bm{p}},x,\xi;k,\gamma,\delta\right)=\mathrm{softmax}\bigl{(}\log({\bm{p}})+\delta\cdot f_{\text{hash}}^{\text{KGW}}\left(x_{\mathrm{len}(x)-k+1},\ldots,x_{\mathrm{len}(x)};\xi,\gamma,|\mathcal{V}|\right)\bigr{)}$ (5)
where $f_{\text{hash}}^{\text{KGW}}$ is a pseudorandom hash function parameterized by key $\xi$ that hashes the previous $k$ tokens in the sequence and returns $g\in\{0,1\}^{|\mathcal{V}|}$, which contains $\gamma\cdot|\mathcal{V}|$ ones and $(1-\gamma)\cdot|\mathcal{V}|$ zeros, encoding the green list. For $k>1$, to hash multiple tokens, we use the Additive-LeftHash scheme (Kirchenbauer et al., 2023b), which adds together the $k$ token IDs. For $k=0$, $f_{\text{hash}}^{\text{KGW}}$ returns a fixed green list $g$ regardless of the previous tokens (Zhao et al., 2023a).
The KGW watermark detection function is
$f_{\text{d}}^{\text{KGW}}\left(x,\xi;\gamma\right)=1-F_{B}\underbrace{\left(\sum\nolimits_{t=k+1}^{\mathrm{len}(x)}f_{\text{hash}}^{\text{KGW}}\left(x_{t-k},\ldots,x_{t-1};\xi,\gamma,|\mathcal{V}|\right)_{x_{t}}\right)}_{\text{number of green list tokens in text $x$}}$ (6)
where $F_{B}$ is the cumulative distribution function (CDF) for binomial distributed random variable $B\sim\mathrm{Bin}(\mathrm{len}(x)-k,\gamma)$. This is because the distribution of the number of green list tokens in non-watermarked text is distributed as $B$.
A.2 Aar
Formally, the Aar (Aaronson, 2023) decoding-based watermarking strategy can be defined as
where $f_{\text{hash}}^{\text{Aar}}$ is a pseudorandom hash function parameterized by key $\xi$ that hashes the previous $k$ tokens and returns $r\in[0,1]^{|\mathcal{V}|}$ with entries uniformly distributed across $[0,1]$, assigning a score to each token in the vocabulary. To hash multiple tokens, the $k$ token IDs are added together. $\mathrm{onehot}(i,|\mathcal{V}|)$ returns a $|\mathcal{V}|$-dimensional probability vector with $1$ at index $i$ and 00 everywhere else, representing deterministic selection of the next token.
The Aar watermark detection function is
where $F_{G}$ is the CDF for gamma distributed random variable $G\sim\mathrm{Gamma}(\mathrm{len}(x)-k,1)$. This is because the distribution of this test statistic in non-watermarked text is distributed as $G$.
A.3 KTH
In the KTH (Kuditipudi et al., 2023) watermarking strategy, the key $\xi$ is a sequence $\xi^{(1)},\ldots,\xi^{(m)}$ where each $\xi^{(j)}\in[0,1]^{|\mathcal{V}|}$ with entries uniformly distributed across $[0,1]$. $m$ should be longer than the maximum generation length. Then, the KTH decoding-based watermarking strategy can be defined as
where $\mathrm{onehot}(i,|\mathcal{V}|)$ returns a $|\mathcal{V}|$-dimensional probability vector with $1$ at index $i$ and 00 everywhere else, representing deterministic selection of the next token.
To allow different generations from the same prompt, before generating each sequence, $\xi$ can be shifted by some random $\tau$, i.e., $\xi^{\prime}=\xi^{(1+\tau\bmod m)},\ldots,\xi^{(m+\tau\bmod m)}$, then $\xi^{\prime}$ is used in $f_{\text{w}}^{\text{KTH}}$. To study the impact of these shifts on learnability, we introduce a hyperparameter $s\in[1,m]$ for how many shift values $\tau$ are possible. We space the $s$ shifts evenly across $[1,m]$, e.g., the set of possible shifts $\tau$ is $\{i\cdot\lfloor m/s\rfloor:0\leq i<s\}$. Increasing $s$ expands the range of possible model generations.
At detection time, to be robust to text edits and shifts, the test statistic quantifies how well a text $x$ can be aligned with the key sequence $\xi$. More specifically, the test statistic computes a minimum Levenshtein distance using the alignment cost $d(x,\xi)=\sum_{t=1}^{\mathrm{len}(x)}\log(1-\xi^{(t)}_{x_{t}})$. See Definition 5 (simple Levenshtein cost) and Equation 6 (alignment cost for exponential minimum sampling) in Kuditipudi et al. (2023). As in Kuditipudi et al. (2023), we set the insertion/deletion penalty $\gamma=0$. A lower (more negative) test statistic indicates stronger watermark signal. To compute p-values, the observed test statistic is compared to a reference distribution of test statistics of non-watermarked texts. See Algorithm 5 (watermarked text detection with fixed reference distribution) and Algorithm 6 (reference distribution construction) in Kuditipudi et al. (2023). Letting $T$ be the number of samples in the reference distribution, the p-values computing using this method are lower bounded by $\frac{1}{T+1}$.
Appendix B Watermark distillation training details
B.1 Logit-based watermark distillation training details
@@
We train Llama 2 7B using a subset of OpenWebText for 5,000 steps with a batch size of 64 sequences, sequence length of 512 tokens, maximal learning rate of 1e-5, and cosine learning rate decay with a linear warmup for the first 500 steps, and the AdamW optimizer (Kingma & Ba, 2015; Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Each training run took approximately 6 hours on 4 NVIDIA A100 80GB GPUs.
B.2 Sampling-based watermark distillation training details
@@
After generating the watermarked samples, we fine-tune Llama 2 7B on the watermarked samples for 1 epoch of 5,000 steps with a batch size of 128 sequences, sequence length of 256 tokens, maximal learning rate of 1e-5, cosine learning rate decay with a linear warmup for the first 500 steps, and the AdamW optimizer (Kingma & Ba, 2015; Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Each training run took approximately 4 hours on 4 NVIDIA A100 80GB GPUs.
Appendix C Pythia sampling-based watermark distillation experiments
We run additional sampling-based watermark distillation experiments using Llama 2 7B as the teacher model and Pythia 1.4B (Biderman et al., 2023) as the student model. Note that Llama 2 and Pythia have different tokenizers, so sampling-based distillation is necessary for this setting.
Training. First, we use Llama 2 7B with a decoding-based watermarking strategy to generate 640,000 watermarked samples of length 256 tokens, prompted with 50-token prefixes from OpenWebText. Then, we fine-tune Pythia 1.4B on the watermarked samples for 1 epoch, roughly 8,000 steps, with a batch size of 64, sequence length of 256 tokens, [^11] Note that Llama 2 and Pythia tokenize sequences differently. Pythia has a larger vocabulary size and tends to tokenize sequences into fewer tokens compared to Llama 2. maximal learning rate of 1e-5, cosine learning rate decay with a linear warmup for the first 500 steps, and the AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Each training run took approximately 3 hours on 1 NVIDIA A100 80GB GPU.
Evaluation. We use the same evaluation procedure and metrics as described in §4.3. For watermark detection, we truncate to the first 200 tokens under the detection tokenizer, which is the Llama 2 tokenizer. Perplexity is still computed using Llama 2 13B. To compute seq-rep-3, we use the Pythia tokenizer.
For each decoding-based watermarking strategy $f_{\text{w}}$, we compare the teacher Llama 2 7B model with $f_{\text{w}}$ applied (denoted by “Decoding (T)”) against the sampling-based distilled student Pythia 1.4B model (denoted by “Sampling”). As a baseline for generation quality, we use the base Pythia 1.4B model with no watermarking or distillation (denoted by “Base student”). Since the teacher Llama 2 7B model is larger and more powerful than the student Pythia 1.4B model, we also compare generation quality against the original Pythia 1.4B model with $f_{\text{w}}$ (denoted by “Decoding (S)”) to control for model size. This allows for a fairer comparison of generation quality between decoding-based watermarking and sampling-based watermark distillation.
Results. Table 3 contains results for the Pythia 1.4B sampling-based watermark distillation experiments. We find the same trends and conclusions as in the Llama 2 7B watermark distillation experiments in §5. Sampling-based distillation using Pythia 1.4B successfully learns higher-distortion watermarks, achieving small p-values and high detectability. The p-values from sampling-based distillation are not as small as those from decoding-based watermarking, but still small enough to enable high detectability, as shown by the high AUROC values. Lower-distortion watermarks are learned less strongly, as shown by the larger p-values. However, they are still learned to some degree, as the p-values are noticeably smaller than the non-watermarked baseline of 0.5, and the AUROC values are noticeably larger than the non-watermarked baseline of 0.5.
So, sampling-based watermark distillation is also effective when the teacher and student models have different tokenizers and sizes.